OpenStudy (butterflyhope):

Urgent *will medal*: How do you do the problem "18 is what percent of 9"? I tried doing it but I kept getting the wrong answer... Thanks c:

hartnn (hartnn):

what have you tried ?

OpenStudy (kiakat):

what's 100% of 9?

OpenStudy (anonymous):

@Butterfly, I helped you with this earlier I thought!

OpenStudy (anonymous):

Oh this is different hehe

hartnn (hartnn):

in math, when unsure, take it as 'x' 18 is x% of 9

hartnn (hartnn):

x% of 9 is (x/100)*9

hartnn (hartnn):

18 = (x/100)*9 get x :)

OpenStudy (anonymous):

.18*9=1.62

OpenStudy (dan815):

18 is 2*9

hartnn (hartnn):

Lizzy, thats for 18% of 9

OpenStudy (kiakat):

@hartnn that's pretty confusing actually...

OpenStudy (dan815):

2 is how many percent?

OpenStudy (dan815):

200

hartnn (hartnn):

agreed, this math way is quite indirect

OpenStudy (anonymous):

right

OpenStudy (dan815):

lets try another example

hartnn (hartnn):

simpler thing to do is 9 is 100% of 9 18 will be 200% of 9 :)

OpenStudy (dan815):

25 is how much % of 5?

OpenStudy (kiakat):

thats what i was trying to say haha

OpenStudy (dan815):

5*5 =25 so % = percent = per100 what over 100 = 5? 500/100 = 5 so 500%

OpenStudy (aripotta):

idk i like to use cross multiplication. gives a visual of what's going on.

OpenStudy (aripotta):

18(100) = 9x and then just solve for x

OpenStudy (dan815):

all % is asking is to represent your factor as a factor with denominator 100

hartnn (hartnn):

butterfly, if you don't know this yet, make a note : \(\Large x\% = \dfrac{x}{100}\)

OpenStudy (butterflyhope):

I got 200%

hartnn (hartnn):

thats \(\huge \checkmark \) :)

OpenStudy (butterflyhope):

Thanks c:

hartnn (hartnn):

welcome :3